OpenStudy (venomblast):

A volcanic block is ejected at an angle of 45 degrees from Mount Fuji during a volcanic eruption. It lands at the foot of the volcano at an horizontal distance of 9.4 km. The height of Mount Fuji is 3.3 km. What is the block's initial speed? Do I use Pythagorean theorem to find the hypotenuse? Is my distance known?

OpenStudy (venomblast):

I got the formula x=x(initial)=v(inital x)*t+1/2at^2

OpenStudy (anonymous):

Nope you would use this formula \[R=\frac{ v _{i}^2\sin2\theta }{ g }\] where R is the horizontal distance it traveled vi is intial velocity theta is the angle g is the acceleration of gravity

OpenStudy (anonymous):

Your formula is incorrect because it is a 2 dimensional motion (x and y component changes) You can ignore the height of the volcano

OpenStudy (anonymous):

But to find R convert the horizontal distance of km to meters

OpenStudy (venomblast):

The think is I never learned that formula/ This is a calc base physic. So I need to to get the velocity by figuring out what I have. I drew a triangle with the angle I get \[V _{0x}= v _{0}*\cos(45)\] \[V _{0y}= v _{0}*\sin(45)\]

OpenStudy (venomblast):

besides I don't know what's R and the initial velocity

OpenStudy (anonymous):

You would need initial velocity to use that formula of yours. It would be the x and y component of the initial velocity. Where were you given "that" in the question?

OpenStudy (venomblast):

there is not a initial velocity. This is suppose to make you find it. So I was thinking. when t(time) is 1/2 does that equal 0 for the y component since it is tangible half way making no velocity when time reaches 1 half of it's destination?

OpenStudy (anonymous):

\[v _{xi}=v _{i}\cos \theta \] \[v _{yi}=v _{i}\sin \theta \]

OpenStudy (venomblast):

I did that. But what V initial?"

OpenStudy (anonymous):

Thats what I said. You mentioned the formula I wrote and I told you that it wasnt needed.

OpenStudy (venomblast):

But I dont think my professor wants to see that since I have not learned it yet. He want's me to use what I have which is the distance formula and velocity/

OpenStudy (venomblast):

so the v( y initial) is irrelevant?

OpenStudy (anonymous):

You sure? Check your notes or online lessons?

OpenStudy (venomblast):

yes. for a thrown object, a=g v=gt+v x=1/2 t^2 + v(initial)

OpenStudy (anonymous):

Its a projectile motion.  It is 2D dimensional